OpenStudy (anonymous):

The first quartile of a data set is 52, and the third quartile is 72. Which of these values in the data set is an outlier? A. 101 B. 41 C. 21 D. 81

OpenStudy (anonymous):

@wio @mathmale

OpenStudy (anonymous):

@thushananth01

OpenStudy (anonymous):

hello wio

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

Well, the middle (2nd quartile) is going to be: \[ \frac{52+72}{2} = \frac{124}{2} = 62 \]So which number if furthest away from \(62\)?

OpenStudy (anonymous):

what is all this [/{}/] stuff

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

but A?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

@Photon336

OpenStudy (anonymous):

No, 101 is not the furthest away

OpenStudy (anonymous):

wut do u mean furthest away

OpenStudy (anonymous):

21 is the furthest away from the middle

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

im still kinda confused. can u help with another

OpenStudy (anonymous):

I think it is C

OpenStudy (anonymous):

It's the best candidate, being furthest away